OpenStudy (anonymous):

How do I factor 100x^2 completely?

OpenStudy (anonymous):

Maybe you can use that 100=10^2

OpenStudy (anonymous):

What can we no next?

OpenStudy (anonymous):

*do next

OpenStudy (anonymous):

@Rheghmë Any clue about how can we use this? 100=10^2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the question that goes with this problem I posted a minute ago. please look at that instead if ur still interested in helping me.

OpenStudy (anonymous):

100x^2=10^2*x^2 And now?